OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 5x = y + 6 2x - 3y = 4 What is the value of the y-coordinate?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What does the word "substitution" mean to you?

OpenStudy (anonymous):

from the first equation y=5x-6 insert this into the second equation

OpenStudy (anonymous):

to subtract from another number

OpenStudy (skullpatrol):

No, that would be "subtraction" this word is "substitution."

OpenStudy (skullpatrol):

Substitution is the process of substituting one thing for another.

OpenStudy (skullpatrol):

Like a substitute teacher is a teacher who is used as replacement for your regular teacher, right?

OpenStudy (anonymous):

ah alright

OpenStudy (skullpatrol):

The substitution principle states an expression may be replaced by another expression that has the same value. Like 4 + 2 can be replaced by 6 because 4 +2 and 6 have the same value, right? Here 4 + 2 and 6 are called numerical expressions.

OpenStudy (anonymous):

Ok

OpenStudy (skullpatrol):

This principle can be generalized to variable expressions as well, like 4x + 2x can be replaced by 6x because 4x + 2x = (4 + 2)x = 6x, using the distributive property.

OpenStudy (anonymous):

Ok im stating to understand

OpenStudy (skullpatrol):

Good :-)

OpenStudy (anonymous):

would that mean the answer would be like 14/13?

OpenStudy (skullpatrol):

That is the answer for the x value. The question asks for the y value. Solve the following system of equations by the substitution method. 5x = y + 6 2x - 3y = 4 What is the value of the y-coordinate?

OpenStudy (skullpatrol):

So use the principle of substitution and replace an x with 14/13 and solve for y.

OpenStudy (anonymous):

would the answer be -16?

OpenStudy (skullpatrol):

You can check to see if your answer is correct by substitution into the the original equation to see if you get a true statement.

OpenStudy (skullpatrol):

I get y = -8/13.